### v1

# BMRetriever: Tuning Large Language Models as Better Biomedical Text Retrievers

Abstract
Developing effective biomedical retrieval models is important for excelling at knowledge-intensive biomedical tasks but still challenging due to the deficiency of sufficient publicly annotated biomedical data and computational resources. We present BMRetriever, a series of dense retrievers for enhancing biomedical retrieval via unsupervised pre-training on large biomedical corpora, followed by instruction fine-tuning on a combination of labeled datasets and synthetic pairs. Experiments on 5 biomedical tasks across 11 datasets verify BMRetriever’s efficacy on various biomedical applications. BMRetriever also exhibits strong parameter efficiency, with the 410M variant outperforming baselines up to 11.7 times larger, and the 2B variant matching the performance of models with over 5B parameters. The training data and model checkpoints are released at https://huggingface.co/BMRetriever to ensure transparency, reproducibility, and application to new domains.
BMRetriever: Tuning Large Language Models as Better
Biomedical Text Retrievers
Ran Xuthanks:  Equal contribution., Wenqi Shi11footnotemark: 1, Yue Yu11footnotemark: 1, Yuchen Zhuang, Yanqiao Zhu May D. Wang, Joyce C. Ho, Chao Zhang, Carl Yang Emory University     Georgia Tech     UCLA {ran.xu,joyce.c.ho,j.carlyang}@emory.edu, yzhu@cs.ucla.edu {wqshi,yueyu,yczhuang,maywang,chaozhang}@gatech.edu
1 Introduction
In the field of biomedicine, the ability to effectively retrieve knowledge from external corpora is crucial for large language models (LLMs) to excel at biomedical NLP tasks (Lewis et al., 2020; Zhang et al., 2024b; Asai et al., 2024). By tapping into up-to-date domain knowledge, retrieval-augmented LLMs have demonstrated promising results in various biomedical downstream applications, including knowledge discovery (Frisoni et al., 2022), question answering (Wang et al., 2023, 2024a), and clinical decision-making (Naik et al., 2022; Shi et al., 2023; Xu et al., 2024).
Figure 1: The average performance of BMRetriever on 5 popular biomedical search tasks compared to baselines with different parameters. X-axis in log scale.
Several works have designed specialized retrieval models for biomedical domains  Mohan et al. (2017); Liu et al. (2021); Jin et al. (2023); Luo et al. (2022a); Singh et al. (2023); Zhang et al. (2023). However, these models are typically built upon BERT-series models, which have limited representative power. Besides, they often rely on proprietary data (e.g., private search logs or patient records), making it challenging to scale them up to accommodate larger models effectively due to privacy concerns. While recent studies in the general domain have improved neural retrieval models via scaling up model size Ni et al. (2022); Muennighoff (2022); Wang et al. (2024b) and training data (Izacard et al., 2022; Wang et al., 2022b; Lin et al., 2023), adapting such models to the biomedical domain may lead to suboptimal performance due to the distribution shift issue (Thakur et al., 2021). Developing large-scale retrieval models dedicated to the biomedical domain without requiring massive proprietary datasets remains crucial yet challenging.
In this work, we propose BMRetriever, a series of dense text retrievers at various scales using LLMs as backbones to improve biomedical retrieval performance. Firstly, we inject biomedical knowledge into BMRetriever by unsupervised contrastive pre-training on a large-scale unlabeled biomedical corpora, which comprises an extensive and diverse collection of data, with rich biomedical background knowledge invaluable for domain-specific understanding (Lála et al., 2023; Xiong et al., 2024). Besides, unlabeled corpora are readily accessible, overcoming the bottleneck of scarce annotated data that often plagues specialized domains. Pre-training on them allows us to adapt our models to the biomedical domain, equipping them with necessary linguistic patterns and terminology.
To further boost the embedding quality and align the retriever with downstream applications, we conduct instruction fine-tuning with high-quality labeled datasets. Specifically, we gather various public human-annotated biomedical retrieval tasks, such as medical question-answering (QA) and dialogue pairs, and create instructions for each to improve BMRetriever with task-specific understanding. Given the relatively small sample size and limited task types in public biomedical datasets, we further leverage the powerful GPT models to generate additional synthetic retrieval tasks under various scenarios with query and passage pairs to augment training samples and diversify instructions. This allows the model to acquire a comprehensive understanding of biomedical retrieval tasks and facilitates its generalization across various downstream tasks and input formats.
We conduct extensive experiments across five tasks on eleven biomedical datasets to demonstrate the strong performance of BMRetriever. As shown in Figure 1, BMRetriever outperforms existing dense retrievers with orders of magnitude more parameters: with 410M parameters, it surpasses the performance of GTR-4.8B (Ni et al., 2022) and SGPT-2.7B (Muennighoff, 2022), which have 7×7$\times$7 × more parameters. At the 7B scale, BMRetriever outperforms the recently proposed E5-Mistral (Wang et al., 2024b), which uses extra-large batch-size and nonpublic data mixture. In addition, BMRetriever presents a lightweight yet high-performing domain adaptation solution, with its 1B variant achieving more than 98% performance of E5-Mistral using only 14.3% of parameters. Our contribution can be summarized as follows:
  • We develop a family of BMRetriever models ranging from 410M to 7B parameters, achieving efficient scaling via a two-stage framework to improve biomedical text retrieval performance.
  • We assess BMRetriever’s efficacy with an extensive evaluation against 18 baselines on 5 tasks across 11 biomedical datasets. Results demonstrate BMRetriever’s parameter efficiency yet strong domain adaptation capabilities, achievable within academic computational budgets.
  • BMRetriever ensures transparency, reproducibility, and potential generalization to additional domain-specific adaptations by providing a detailed training recipe with public datasets and accessible model checkpoints.
2 Related Work
Earlier research explores various approaches for learning representations suitable for text retrieval (Deerwester et al., 1990; Huang et al., 2013). More recently, several studies introduce dual-encoder architectures based on BERT for dense retrieval Karpukhin et al. (2020); Xiong et al. (2021); Qu et al. (2021); Gao and Callan (2022); Izacard et al. (2022); Xiao et al. (2022). With the advent of LLMs with billions of parameters, several studies attempt to scale up model size (Ni et al., 2022; Neelakantan et al., 2022; Muennighoff, 2022; Springer et al., 2024), often fine-tuned on multi-task instruction data (Asai et al., 2023; Su et al., 2023; Wang et al., 2024b; Muennighoff et al., 2024). However, the benefit of scaling up is more pronounced for general domain datasets where massive annotated data are available.
To design effective retrievers for specialized domains, several works propose continuously pre-train the retrieval model on domain-specific corpora (Yu et al., 2022; Zhang et al., 2023) or fine-tuning the model on proprietary search datasets (Mohan et al., 2017; Jin et al., 2023). On the other hand, synthetic data has also been used to improve the generalization ability of dense retrieval model (Ma et al., 2021; Wang et al., 2022a; Jiang et al., 2023; Wang et al., 2024b; Zhang et al., 2024a). Despite these advancements, how to combine public, open data to formulate a dataset curation recipe for adapting LLMs as high-performing biomedical retrievers remains unresolved. Our method efficiently integrates diverse supervision signals for biomedical retrieval model training, which achieves better performance than baselines trained with more data.
3 Method
Figure 2: The overview of the two-stage pre-training framework in BMRetriever. Stage I performs unsupervised contrastive pre-training on large-scale biomedical query-passage pairs, while Stage II conducts instruction fine-tuning using diverse labeled data, including synthetic examples generated by LLMs, to adapt BMRetriever to various biomedical downstream tasks.
Parameters 410M 1B 2B 7B
Backbone Pythia Biderman et al. (2023) Pythia Biderman et al. (2023) Gemma Team et al. (2024) BioMistral Labrak et al. (2024)
Model Layers 24 16 18 32
Embedding Dim. 1024 2048 2048 4096
Table 1: An overview of BMRetriever.
BMRetriever leverages the pre-trained autoregressive transformer as the backbone, taking advantage of the availability of various model sizes within this model family. This flexibility allows us to scale up the retrieval model. Specifically, we utilize the publicly available autoregressive transformers with 410M, 1B, 2B, and 7B parameters (Biderman et al., 2023; Team et al., 2024; Labrak et al., 2024). Our model details are illustrated in Table 1.
3.1 Background of Dense Text Retrieval
In dense retrieval (Lee et al., 2019; Karpukhin et al., 2020), the language model $\mathbf{E}$ is used to represent queries and passages in dense embeddings. Denote the query $q$ and passage $p$ with the corresponding task instruction $I_{q}$ and $I_{p}$ [^1] The instruction format is in Appendix B., the embedding is calculated as $\bm{e}_{q}=\mathbf{E}\left(I_{q}\oplus q\right)$, $\bm{e}_{p}=\mathbf{E}\left(I_{p}\oplus p\right)$. The relevance score $\operatorname{sim}(q,p)$ is calculated with the dot product between query and passage embeddings:
$\operatorname{sim}(q,p)=\bm{e}_{q}^{\top}\bm{e}_{p}.$ (1)
In this work, where autoregressive LLMs are used for $\mathbf{E}$, an <EOS> token is appended to the end of the query and passage. The embedding of the <EOS> token from the final layer of LLM is used as the representation for both queries and passages.
To effectively adapt BMRetriever to the biomedical domain, a two-stage training procedure is proposed (see Figure 2): (1) an unsupervised contrastive pre-training stage (§ 3.2) using silver query-passage pairs from extensive biomedical corpora, and (2) a fine-tuning stage (§ 3.3) using gold labeled data from various tasks. The details of the two stages will be introduced in the following sections.
3.2 Unsupervised Contrastive Pre-training
Pre-training Corpus Collection. To provide BMRetriever with an initial understanding of biomedical contexts, we collect a diverse range of publicly available biomedical corpora, including biomedical publications (Chen et al., 2021; Xiong et al., 2024; Lo et al., 2020), medical textbooks (Jin et al., 2021), as well as general-domain web corpus (Bajaj et al., 2016), as detailed in Table 8.
Contrastive Pre-training. We construct positive and negative query-passage pairs from raw unlabeled corpora to facilitate contrastive pre-training of the retrieval model. For positive pairs, we employ two strategies: (1) for corpora with titles, we treat the title as the query and the corresponding abstract as the passage; (2) for untitled corpora, we randomly sample two disjoint passages from documents, using one as the query and the other as the passage (Izacard et al., 2022). To obtain negative pairs, we sample in-batch negatives (Gillick et al., 2019) where the passages from other pairs in the same batch serve as negative examples. With the collected pairs, we employ contrastive learning to distinguish the relevant query-passage pairs from the irrelevant ones. For each mini-batch $\mathcal{B}$, we leverage the InfoNCE loss as the pre-training objective to rank the positive text pairs $\{(q_{i},p_{i})\}_{i=1}^{n}$ higher than in-batch negative passages $\{p_{ij}^{-}\}_{j=1}^{N}$:
$\mathcal{L}_{\text{cpt}}=-\log\frac{{e}^{\operatorname{sim}\left({q_{i}},p_{i}\right)/\tau}}{\sum_{j\in\mathcal{B}}{e}^{\operatorname{sim}\left(q_{i},p_{j}\right)/\tau}}.$ (2)
Contrastive pre-training improves the quality of representations by better aligning similar text sequences while ensuring the uniformity of unrelated text sequences, which helps adapt the retrieval model to biomedical domains (Gururangan et al., 2020; Yu et al., 2022; Luo et al., 2022b).
3.3 Supervised Instruction Fine-tuning
To further enhance the model’s specialized domain knowledge and align the model with downstream application tasks, we conduct instruction fine-tuning, which integrates a diverse collection of retrieval tasks into the instruction tuning blend. We present a detailed procedure below.
Task Scale # PT Pairs # FT Pairs Standard IR Sent. Sim. Avg. Retr. Avg. All
Model NFCorpus SciFact SciDocs Trec-COVID BIOSSES
Sparse Retrieval
BM25 (Robertson et al., 2009) 0.325 0.665 0.158 0.656 0.451
Base Size (< 1B)
Contriever (Izacard et al., 2022) 110M 1B 500K 0.328 0.677 0.165 0.596 0.833 0.442 0.520
Dragon (Lin et al., 2023) 110M 28.5M 0.339 0.679 0.159 0.759 0.819 0.484 0.551
SPECTER 2.0 (Singh et al., 2023) 110M 3.3M 0.228 0.671 0.584
SciMult (Zhang et al., 2023) 110M 5.5M 0.308 0.707 0.712
COCO-DR (Yu et al., 2022) 110M 15M 500K 0.355 0.709 0.160 0.789 0.829 0.503 0.567
SGPT-125M (Muennighoff, 2022) 125M unknown 500K 0.228 0.569 0.122 0.703 0.752 0.406 0.475
MedCPT (Jin et al., 2023) 220M 255M 0.340 0.724 0.123 0.697 0.837 0.471 0.544
GTR-L (Ni et al., 2022) 335M 2B 662K 0.329 0.639 0.158 0.557 0.849 0.421 0.506
InstructOR-L (Su et al., 2023) 335M 1.24M 0.341 0.643 0.186 0.581 0.844 0.438 0.519
E5-Large-v2 (Wang et al., 2022b) 335M 270M 1M 0.371 0.726 0.201 0.665 0.836 0.491 0.560
BGE-Large∗‡ (Chen et al., 2024) 335M 1.2B 1.62M 0.345 0.723 0.222 0.753 0.804 0.511 0.569
BMRetriever-410M 410M 10M 1.4M 0.321 0.711 0.167 0.831 0.840 0.508 0.574
Large Size (1B - 5B)
InstructOR-XL (Su et al., 2023) 1.5B 1.24M 0.360 0.646 0.174 0.713 0.842 0.473 0.547
GTR-XL (Ni et al., 2022) 1.2B 2B 662K 0.343 0.635 0.159 0.584 0.789 0.430 0.502
GTR-XXL (Ni et al., 2022) 4.8B 2B 662K 0.342 0.662 0.161 0.501 0.819 0.417 0.497
SGPT-1.3B (Muennighoff, 2022) 1.3B unknown 500K 0.320 0.682 0.162 0.730 0.830 0.473 0.545
SGPT-2.7B (Muennighoff, 2022) 2.7B unknown 500K 0.339 0.701 0.166 0.752 0.848 0.489 0.561
BMRetriever-1B 1B 10M 1.4M 0.344 0.760 0.180 0.840 0.858 0.531 0.596
BMRetriever-2B 2B 10M 1.4M 0.351 0.760 0.199 0.863 0.828 0.543 0.600
XL Size (> 5B)
SGPT-5.8B (Muennighoff, 2022) 5.8B unknown 500K 0.362 0.747 0.199 0.849 0.863 0.539 0.604
LLaRA (Li et al., 2023a) 7B 21M 500K 0.372 0.757 0.172 0.853 0.539
RepLLaMA (Ma et al., 2023) 7B 500K 0.378 0.756 0.181 0.847 0.541
Grit-LM (Muennighoff et al., 2024) 7B 2.2M 0.408 0.792 0.244 0.748 0.864 0.548 0.611
LLM2Vec (BehnamGhader et al., 2024) 7B 1.2M 1.5M 0.393 0.788 0.225 0.776 0.852 0.545 0.606
E5-Mistral (Wang et al., 2024b) 7B 1.8M 0.386 0.764 0.162 0.872 0.855 0.546 0.608
CPT-text-XL (Neelakantan et al., 2022) 175B unknown unknown 0.407 0.754 0.649
BMRetriever-7B 7B 10M 1.4M 0.364 0.778 0.201 0.861 0.847 0.551 0.610
Table 2: Main experiments on biomedical text representation tasks in various scales. Bold and underline indicate the best and second best results on average performance over the four retrieval tasks, and over all five tasks. * denotes concurrent works (for reference only). $\dagger$ uses reranker distillation. $\ddagger$ employs hybrid retrieval. We highlight the biomedical or scientific domain-specific retrieval models. Notations are consistent across tables. “PT”, “FT”, and “Sent. Sim.” denote “Pre-training”, “Fine-tuning”, and “Sentence Similarity”, respectively.
Instruction Fine-tuning Dataset. To incorporate the model with a wide range of biomedical downstream tasks, we leverage a series of biomedical tasks with varying granularity, including both sentence-level medical natural language inference (MedNLI) (Shivade, 2017), medical question pairs (McCreery et al., 2020), and passage-level biomedical QA tasks, including MedMCQA (Ben Abacha et al., 2019), StackExchange (Team, 2021), and medical dialogues (Li et al., 2023b). Besides, we also include several general-domain retrieval datasets, including MS MARCO Bajaj et al. (2016), NQ (Kwiatkowski et al., 2019), Fever (Thorne et al., 2018), ELI5 (Fan et al., 2019), and NLI (Bowman et al., 2015), to enhance the model’s ability for relevance estimation. The instruction format and data conversion details are exhibited in Appendix B.
Synthetic Data Augmentation with LLMs. To supplement the limited task types and relatively small sample sizes in labeled biomedical datasets, we employ a data augmentation approach to generate synthetic query and passage pairs. Two approaches are utilized for this generation process.
We leverage GPT-3.5 (gpt-3.5-turbo-1106) for instance-level augmentation to enrich (query, passage) pairs resembling standard biomedical information retrieval (IR) formats. Given a passage from PubMed and Meadow used in contrastive pre-training, we prompt GPT-3.5 to generate a relevant query based on the passage context. This allows the model to better capture the relevance within biomedical contexts for effective retrieval.
Beyond relevance signals, task generalization is also crucial for building a general retriever, as user intent and input formats vary while public data captures only a fraction of tasks. To address this, we perform task-level augmentation, which involves prompting GPT-4 (gpt-4-turbo-1106) to conceptualize a diverse list of potential scenarios for biomedical retrieval tasks (Wang et al., 2024b). Subsequently, we prompt GPT-4 again to generate examples for each scenario, including a query, a relevant (positive) passage, and a challenging irrelevant (hard negative) passage. This approach allows us to enhance the diversity of instructions.
Hard Negative Mining and Data Filter. In both labeled instruction fine-tuning datasets and data-label synthetic datasets, positive pairs are available, while negative examples are missing. To obtain the negatives, we randomly select 1 passage from the top 100 passages retrieved by E5-base (Wang et al., 2022b) when using the given query to search the entire corpus of the corresponding dataset. As the generated synthetic data can be noisy, consistency filtering is adopted to filter low-quality pairs (Alberti et al., 2019; Dai et al., 2023), where for each synthetic (query $q$, passage $p$) pair, we use the E5-base to predict the most relevant passages for ${q}$. We only retain $q$ when $p$ occurs among the top three retrieved passages.
Fine-tuning Objectives. After constructing positive and negative text pairings $\{(q_{i},p_{i}^{+},p_{i}^{-})\}_{i=1}^{M}$ where $p_{i}^{+}$ and $p_{i}^{-}$ stands for the positive passage and the hard negative, respectively, we employ the InfoNCE loss function for each minibatch $\mathcal{B}$ as
$\mathcal{L}_{\text{ft}}=\frac{e^{\operatorname{sim}(q_{i},p_{i}^{+})/\tau}}{\sum_{j\in\mathcal{B}}e^{\operatorname{sim}(q_{i},p_{j}^{+})/\tau}+e^{\operatorname{sim}(q_{i},p_{j}^{-})/\tau}},$ (3)
where both in-batch negatives and hard negatives are utilized to further improve model training.
4 Experimental Results
4.1 Experimental Setups
Tasks and Datasets. We conduct experiments on eleven datasets across five biomedical retrieval-oriented tasks, including (1) IR, (2) sentence similarity (STS), (3) QA, (4) entity linking, and (5) paper recommendation. There is no overlap between the training and test pairs. Task and dataset details are available in Appendix B.
Baselines. We compare to sparse retrieval models BM25 (Robertson et al., 2009) and open-source dense retrieval models with varying model sizes: Contriever (Izacard et al., 2022), Dragon (Lin et al., 2023), SciMult (Zhang et al., 2023), SPECTER 2.0 (Singh et al., 2023), COCO-DR (Yu et al., 2022), QExt (Meng et al., 2022), SGPT (Muennighoff, 2022), MedCPT (Jin et al., 2023), GTR (Ni et al., 2022), InstructOR (Su et al., 2023), E5-Large-v2 (Wang et al., 2022b), BGE-Large (Chen et al., 2024), LLaRA (Li et al., 2023a), RepLLaMA (Ma et al., 2023), Grit-LM (Muennighoff et al., 2024), LLM2Vec (BehnamGhader et al., 2024), E5-Mistral (Wang et al., 2024b), and CPT-text (Neelakantan et al., 2022). The details of baselines and parameter sizes are in Appendix C.
Evaluation. We follow existing works (Muennighoff et al., 2023; Thakur et al., 2021) and use nDCG@10 to measure standard IR performance and Spearman correlation for STS based on cosine similarity. To evaluate the retrieval performance of QA, we report Recall@{5,20} and nDCG@20. For entity linking, we report mean reciprocal rank (MRR)@5 and Recall@{1,5}. For paper recommendation, we follow Singh et al. (2023) and report mean average precision (MAP) and nDCG. The implementation details can be found in Appendix D.
Task Question Answering Entity Linking Paper Rec.
Model BioASQ PubMedQA iCliniq DrugBank MeSH RELISH
R@5 R@20 nDCG@20 R@5 R@20 nDCG@20 R@5 R@20 nDCG@20 R@1 R@5 MRR@5 R@1 R@5 MRR@5 MAP nDCG
Base Size (< 1B)
Dragon Lin et al. (2023) 36.2 54.6 49.1 71.8 74.0 72.0 50.6 65.2 47.4 81.0 87.6 83.3 28.2 47.0 34.8 72.6 80.6
MedCPT Jin et al. (2023) 34.7 54.4 45.2 66.3 71.1 60.4 26.8 42.0 24.9 75.1 88.0 80.6 27.7 54.2 37.4 83.6 89.7
E5-Large-v2 Wang et al. (2022b) 36.8 54.0 50.4 71.6 74.2 72.2 57.6 72.0 55.8 81.8 86.5 81.5 32.8 55.0 41.3 84.9 91.0
BMRetriever-410M 39.9 54.2 53.1 73.8 74.6 72.4 60.6 72.8 56.6 81.4 88.2 83.7 31.5 53.8 39.8 85.2 91.2
Large Size (1B - 5B)
InstructOR-XL Su et al. (2023) 29.9 43.2 41.8 70.5 74.0 69.1 64.9 78.1 58.3 75.3 84.2 80.3 33.6 56.2 45.7 84.5 90.6
SGPT-2.7B Muennighoff (2022) 33.9 47.4 47.3 68.3 73.7 63.2 45.0 52.2 41.2 71.9 77.0 62.9 20.2 39.7 28.5 84.9 90.8
BMRetriever-1B 40.4 55.8 53.4 73.6 74.4 72.7 61.1 73.7 56.8 84.7 89.1 86.5 35.5 60.3 48.8 85.2 91.3
BMRetriever-2B 42.5 56.5 55.7 74.0 74.6 73.1 70.0 81.2 65.7 82.6 90.2 85.8 45.6 71.3 59.5 85.4 91.5
XL Size (> 5B)
E5-Mistral Wang et al. (2024b) 39.6 55.4 52.7 72.6 74.2 70.0 56.7 72.2 51.8 78.5 92.2 84.0 47.9 76.2 61.3 85.2 90.8
BMRetriever-7B 43.7 60.2 57.4 74.2 74.6 73.8 68.4 79.7 63.7 84.7 92.8 88.0 49.8 76.5 61.1 86.7 92.2
Table 3: Experiments on retrieval-oriented biomedical NLP applications compared with strongest and fair baselines.
4.2 Results on Text Representation Tasks
Table 2 presents a comprehensive evaluation of the embedding quality on four standard biomedical IR tasks and an additional task focused on biomedical sentence similarity. Across different scales, BMRetriever outperforms the majority of baseline methods, achieving either the highest or second-highest performance in terms of average scores on the four IR tasks, as well as on the combined set of all five tasks. It even outperforms E5-Large-v2 (Wang et al., 2022b) with additional supervision signals and matches BGE-Large’s hybrid retrieval approach combining dense, lexical, and multi-vector retrieval (Chen et al., 2024). Here we focus on scaling up biomedical retrieval models with mixed data types, leaving the combination of BMRetriever with other more complex and larger scale language systems for future work.
A notable aspect of BMRetriever is its efficiency and lightweight nature. Its 410M, 1B, and 2B variants achieve 94.1%, 97.7%, and 98.4% performance using only 5.9%, 14.3%, and 28.6% of 7B variant’s parameters, respectively. Moreover, BMRetriever-410M outperforms all the baselines in large size (1B-5B) with up to 11.7×11.7$\times$11.7 × more parameters, and BMRetriever-2B matches performance with baselines in XL size (> 5B). Remarkably, BMRetriever also provides a reasonable training setup within an academic budget, requiring only 10M pre-training data and 1.5M fine-tuning data, which is significantly less than the data usage in most baselines, such as GTR (Ni et al., 2022) and MedCPT (Jin et al., 2023). Yet, BMRetriever still outperforms these data-intensive methods.
4.3 Results on Retrieval-Oriented Biomedical Applications
Table 3 evaluates BMRetriever’s performance on biomedical downstream applications. The results demonstrate BMRetriever’s efficacy over most baselines across different tasks and datasets, justifying the adaptability of our learned representations to various retrieval-oriented applications.
Furthermore, our proposed BMRetriever exhibits strong generalization capabilities across diverse tasks and input formats, including retrieving long context from short questions (BioASQ, PubMedQA), retrieving long answers from patient questions (iCliniq), retrieving definitions from entity names (DrugBank, MeSH), and retrieving relevant abstracts given an abstract (RELISH). Notably, BMRetriever performs well on unseen tasks, such as entity linking and paper recommendation, verifying its ability to generalize to new tasks unseen in the instruction fine-tuning stage.
Figure 3: Effect of different fine-tuning data on various datasets. “FT” denotes “Fine-tuning”.
4.4 Studies on Contrastive Pre-training
Task Size Standard IR Sent.
Sim. Avg. Avg.
Model NFC. Sci- Sci- Trec- BIO- Retr. All
Fact Docs COVID SSES
Contriever Izacard et al. (2022) 110M 0.328 0.677 0.165 0.274 0.781 0.347 0.434
COCO-DR Yu et al. (2022) 110M 0.243 0.724 0.150 0.483 0.801 0.400 0.480
QExt Meng et al. (2022) 110M 0.303 0.644 0.147 0.535 0.407
E5-Large-v2 Wang et al. (2022b) 335M 0.337 0.723 0.218 0.618 0.822 0.474 0.543
LLM2Vec BehnamGhader et al. (2024) 7B 0.271 0.687 0.153 0.557 0.832 0.417 0.500
BMRetriever 410M 0.306 0.677 0.180 0.802 0.834 0.491 0.560
BMRetriever 1B 0.330 0.744 0.187 0.800 0.833 0.515 0.579
BMRetriever 2B 0.342 0.738 0.198 0.848 0.847 0.531 0.593
BMRetriever 7B 0.355 0.750 0.208 0.833 0.861 0.537 0.601
Table 4: The performance of unsupervised dense retrieval models on biomedical representation tasks. Directly using the backbone model of BMRetriever (before contrastive pre-training) leads to performance $<0.03$ for all datasets, thus we do not report them.
To highlight the effectiveness of our contrastive pre-training approach, we evaluate the performance of unsupervised dense retrieval models that only use unlabeled corpora for pre-training. As shown in Table 4, our model outperforms existing unsupervised baselines and even surpasses many fully supervised models reported in Table 2. The strong unsupervised results have important implications for real-world biomedical applications, where curating large labeled datasets is often prohibitively expensive and time-consuming. Our approach presents an attractive alternative, enabling the development of high-quality retrieval models in a data-efficient manner using only the unlabeled corpus.
(a) Effect of CL strategies
(b) Ablation Studies
Figure 4: Additional results over five tasks in the main experiments. “CL” stands for “Contrastive Learning”.
We further investigate the performance of employing cropping alone as the contrastive pre-training strategy, which entails randomly selecting two passages from the corpus as a positive query-passage pair (Gao and Callan, 2022; Izacard et al., 2022). The results presented in Table 4(a) demonstrate that utilizing cropping as the sole contrastive learning objective yields suboptimal performance.
4.5 Studies on Instruction Fine-tuning
Figure 3 illustrates the impact of different fine-tuning data sources on model performance across various datasets [^2] Removing biomedical data retains the synthetic data.. Among all the utilized data types, synthetic data contributes the most significant performance gain, which can be attributed to its larger volume compared to biomedical data and its coverage of a more diverse range of task types. It is particularly beneficial for NFCorpus, SciFact, and Trec-COVID, as these datasets follow the standard IR format of short queries and long passages, aligning with the format of the synthetic data. Furthermore, synthetic data proves advantageous for the iCliniq dataset, as it potentially includes various retrieval scenarios, such as dialog data. General domain fine-tuning data, consisting of short queries and long passages, generally enhances relevance estimation capabilities, benefiting standard biomedical IR tasks like Trec-COVID. However, it may slightly diminish performance on datasets like iCliniq, which contain conversational patient-provider data that deviates from traditional IR formats, potentially introducing distribution shifts. Biomedical fine-tuning data, on the other hand, demonstrates broad usefulness, particularly for BIOSSES and iCliniq tasks, likely due to its inclusion of sentence similarity and dialog data aligning with downstream task formats. Public fine-tuning data from E5-Mistral (Wang et al., 2024b) and MEDI, used by InstructOR (Su et al., 2023), achieves relatively poor performance, possibly due to their focus on the general domain with limited biomedical knowledge. Additionally, the lack of publicly available synthetic generated data from E5-Mistral (Wang et al., 2024b) may contribute to the suboptimal performance observed.
4.6 Additional Studies
Stage ($\downarrow$) Volume ($\rightarrow$) 10% 50% 100%
Pre-training BMRetriever-410M 0.540 0.554 0.560
BMRetriever-1B 0.564 0.575 0.579
Fine-tuning BMRetriever-410M 0.562 0.571 0.574
BMRetriever-1B 0.590 0.595 0.596
Table 5: Effect of data volume in pre-training and fine-tuning. Pre-training results do not involve subsequent fine-tuning. Fine-tuning results are based on the pre-training checkpoints with full pre-training data.
Case Study 1
Dataset / Task iCliniq / Given a patient query, retrieve its corresponding answer from the community website.
Query My grandmother is having dizziness, vomiting and headache. These problems occur constantly with a gap of five to six months.
BMRetriever From the symptoms that you mentioned it seems your grand mother has migraine which reccurs when she is either exposed to the allergen, which triggers her migraine, or it can also manifest as a seasonal change. How long has she has been having this problem? During her migraine episodes, how long does it last? I will recommend the following medicines :- 1) Tablet Crocin pain relief one tablet twice daily after food. 2) Tablet Ondem (Ondansetron) 4 mg, if there is severe nausea two tablets can be taken at once. Make sure her BP (blood pressure) is under control.
BMRetriever w/o Instruction Please provide details regarding your headache and abuse. Also, provide details about your current symptoms and like how they started and how they progress over time. What are the medications you are taking currently and for how long? Once I receive this information, I will be able to help you.
[…]
E5-Mistral You are suffering from Episodes of headache with nausea, which are not present every day. also having left eye pain and twisting with tingling, imbalance while getting up, tingling in throat and over genitals, coldness over forehead … I would like to suggest to do some routine blood investigation tests and this symptoms seems like vitamin B12 deficiency or hypothyroidism. Are you Vegetarian?
Case Study 2
Dataset / Task SciFact / Given a biomedical claim, retrieve biomedical literature to support or refute the claim.
Query Forkhead 0 (fox0) transcription factors are involved in apoptosis.
[…]
Ablation Studies. We inspect different components of BMRetriever in Figure 4(b), including instruction, pre-training, and fine-tuning [^3] Removing instruction solely eliminates the retrieval guidance, while retaining both pre-training and fine-tuning.. The results indicate that removing any component would hurt the performance. We also observe that pre-training is particularly beneficial for smaller models, as larger models may already possess sufficient capacity to capture domain knowledge.
Effect of Data Volume. Table 5 evaluates the effect of data volume during pre-training and fine-tuning. The results demonstrate the remarkable efficiency of BMRetriever, achieving comparable performance even when trained on substantially less data. Notably, using only 10% of the data, the 1B variant of BMRetriever outperforms all baselines in either the pre-training or fine-tuning stage, while the 410M variant also achieves better performance than most baselines in fine-tuning.
4.7 Case Study
We present two case studies in Table 6 illustrating the quality of retrieved passages from BMRetriever compared to strong baselines. The first example, from the iCliniq dataset, considers a patient query and retrieves the corresponding answer from a community website. In the given example, BMRetriever retrieves a passage directly addressing symptoms like headaches and nausea, recommending medication aligning with the condition. In contrast, the retrieved passage from MedCPT focuses on inner ear problems and vertigo, not covering the vomiting or the specific periodicity of the episodes described in the query. The passage from E5-Mistral talks about symptoms not mentioned by the patient, such as left eye pain and tingling. Besides, we also present the result from BMRetriever without using instructions, which is also imprecise since it mentions abuse, a topic not relevant to the query.
The second example involves retrieving biomedical literature to support or refute a claim about apoptosi. The passage retrieved by BMRetriever specifically mentions that the FoxO family of transcription factors participates in apoptosis. Although the passage retrieved by MedCPT discusses the role of FoxO transcription factors in cell death, it is specific to neuronal cells under ischemic conditions, rather than general apoptosis. Furthermore, both E5-Mistral and BMRetriever without instructions retrieve an irrelevant passage about the role of FoxO1 in regulating regulatory T cells, unrelated to the claim. We further illustrate the cosine similarity distributions of relevant and irrelevant (query, passage) pairs in Appendix F.
5 Conclusion
We present BMRetriever, a series of dense retrieval models designed for knowledge-intensive biomedical NLP tasks with various scales. BMRetriever is pre-trained on a large-scale biomedical corpus and further instruction fine-tuned on diverse, high-quality biomedical tasks. Through extensive experimentation, we have demonstrated that BMRetriever exhibits state-of-the-art performance across a range of biomedical applications. Furthermore, BMRetriever demonstrates impressive parameter efficiency, with its smaller variants achieving 94-98% of the performance of the 7B model using only 6-29% as many parameters, while the 410M version surpasses larger baselines (1B-5B) up to 11.7 times larger. We hope BMRetriever can be incorporated into a broad suite of biomedical tasks to advance biomedical NLP research.
Limitation
Efficiency. One specific caveat for scaling up model size is the increment in the latency overhead. We have reported both the passage indexing speed and retrieval latency in Appendix G, which indicates that our model does not incur much additional time when compared to models with similar size (e.g., BMRetriever-2B v.s. InstructOR-1.5B). One important future work is to explore how to reduce the inference latency and lower the storage cost for text embeddings produced by LLMs.
Cost Estimation. Generating synthetic data using GPT models incurs additional costs. In our work, the total API cost of BMRetriever is less than \$500444As of April 2024., which remains affordable within an academic budget. This cost is significantly lower than recent works (Wang et al., 2024b), which have an estimated cost of more than \$6000.
Other Training Objective and Backbone Models. The primary focus of this work is on building an effective dataset creation recipe to scale up retrieval models for the biomedical domain. However, other biomedical LLMs can also be used as the backbones (Luo et al., 2022b; Bolton et al., 2024; Chen et al., 2023). In our early attempts, we did not observe significant performance gains with some of these backbones (e.g., BioGPT-Large, BioMedLM-2.7B, Meditron-7B), potentially due to the loose connection between the objective of sequence contrastive learning and causal language modeling. Besides, since we utilize standard contrastive learning objectives with hard negative mining without proposing new learning objectives, we do not claim this as our core contribution. Exploring ways to leverage domain-specific LLMs with improved training techniques to enhance performance is an interesting topic for future research.
Ethics Consideration
Misinformation. One specific issue for LLM-generated biomedical text is the potential for misinformation and hallucination (Pal et al., 2023). It is important to note that for the generated queries, the majority are short sentences or phrases without presenting any scientific facts. Regarding the generated (query, passage) pairs, to ensure that our generated synthetic text does not introduce misinformation or hallucination, we randomly selected 200 examples and asked medical students to evaluate the factuality of the generated text. The evaluation results did not reveal any instances of misinformation or hallucination in the selected examples.
Data Contamination. A potential issue is test set contamination (Sainz et al., 2023), where some test examples overlap with the training data. This can be especially problematic for text generated by LLMs, as they are often pre-trained on massive corpora spanning various domains. To address this concern, we follow Wang et al. (2024b) to conduct a string match-based analysis between the test set and our training set, where we do not observe any overlap between the train and test queries. While some of the corpora (e.g., PubMed) are also utlized in the test tasks, this is a standard practice even in zero-shot or few-shot evaluation of retrieval models (Ma et al., 2021; Wang et al., 2022a; Yu et al., 2022; Dai et al., 2023), and it is not considered as contamination.
Appendix A Additional Synthetic Data Augmentation Details
A.1 Prompt Format to Generate Task and Pairs
Listing 1: Prompt format for synthetic retrieval task generation.
Brainstorm a list of potentially useful biomedical text retrieval tasks.
Here are a few examples for your reference:
1. Provided a scientific claim as query, retrieve documents that help verify or refute the claim.
2. Search for documents that answers a FAQ-style query on childrens nutrition.
Please adhere to the following guidelines:
1. Specify what the query is, and what the desired documents are.
2. Each retrieval task should cover a wide range of queries, and should not be too specific.
3. Focus on biomedical related topics.
Your output should always be a python list of strings only, with about 20 elements, and each element corresponds to a distinct retrieval task in one sentence. Do not explain yourself or output anything else. Be creative!
Listing 2: Prompt format for synthetic retrieval examples generation.
You have been assigned a biomedical retrieval task: [task]
Your mission is to write one biomedical text retrieval example for this task in JSON format. The JSON object must contain the following keys:
1. "user_query": a string, a random user search query specified by the retrieval task.
2. "positive_document": a string, a relevant document for the user query.
3. "hard_negative_document": a string, a hard negative document that only appears relevant to the query.
Please adhere to the following guidelines:
1. The "user_query" should be [query_type], [query_length], [clarity], and diverse in topic.
2. All documents should be at least [num_words] words long.
3. Both the query and documents should be in English.
4. Both the query and documents require [difficulty] level education to understand.
Your output must always be a JSON object only, do not explain yourself or output anything else. Be creative!
[task]: The task names generated from the previous step.
[query_type]: Randomly sampled from ["extremely long-tail", "long-tail", "common"].
[query_length]: Randomly sampled from ["less than 5 words", "5-10 words", "at least 10 words"]
[clarity]: Randomly sampled from ["clear", "understandable with some effort", "ambiguous"]
[num_words]: Randomly sampled from ["50 words", "50-100 words", "200 words", "300 words", "400 words"]
[difficulty]: Randomly sampled from ["high school", "college", "PhD"]
A.2 Case Study
Synthetic Data Augmentation Example 1
Synthetic Retrieval Task Retrieve documents that discuss the potential health risks associated with obesity
Query Potential health hazards related to obesity
Positive Passage The escalating obesity epidemic is associated with a range of serious health risks and complications. Obesity is a chronic disease characterized by excessive accumulation and storage of fat in the body, which in turn is associated with a multitude of adverse health effects. These effects include an increased risk of premature death, cardiovascular diseases including hypertension, stroke and coronary artery disease, type 2 diabetes, certain types of cancer, and sleep apnea among others. The psychosocial consequences of obesity are also significant, with reduced quality of life and depression being common. The pathophysiological basis of these complications is complex and involves a myriad of interrelated metabolic, inflammatory, and vascular processes. For instance, the excess adiposity in obesity contributes to systemic inflammation and insulin resistance, which are central to the development of type 2 diabetes and cardiovascular diseases. The increased mechanical load of obesity can also lead to degenerative joint diseases, while the increased metabolic demand can lead to heart failure. Furthermore, the excess adipose tissue can secrete a variety of adipokines and cytokines that can promote carcinogenesis. Thus, the potential health risks associated with obesity are multifold, and the prevention and management of obesity should be a public health priority.
Negative Passage Obesity is increasingly recognized as a major health problem worldwide. While it is associated with a multitude of health risks, the potential benefits of obesity are less well known. Some studies have suggested that obese individuals may have a lower risk of certain conditions such as osteoporosis, due to their higher bone mineral density. Furthermore, some obese individuals seem to be ’metabolically healthy’, with no apparent signs of the metabolic abnormalities typically associated with obesity. These findings, however, do not negate the overall detrimental effects of obesity on health. Rather, they highlight the complexity of obesity as a disease and the need for personalized approaches in its management. It is also important to note that the ’obesity paradox’, where obesity seems to confer survival advantages in certain conditions such as heart failure, remains controversial and is not yet fully understood. Overall, while there may be potential benefits associated with obesity, these are likely outweighed by the substantial health risks, and efforts should be focused on preventing and managing obesity to improve health outcomes.
Synthetic Data Augmentation Example 2
Synthetic Retrieval Task Search for documents that provide information on the latest treatments for autoimmune diseases
Query I am looking for scholarly articles or scientific papers that describe the most recent advancements in therapies for autoimmune diseases, such as rheumatoid arthritis, lupus, celiac disease, or multiple sclerosis.
Positive Passage In recent years, there have been significant advancements in the treatment of autoimmune diseases. One major development is the use of biologics, which are drugs derived from living organisms. Biologics have been successfully used in the treatment of rheumatoid arthritis, lupus, and other autoimmune disorders. They work by targeting specific components of the immune system that cause inflammation and damage. Another promising treatment is stem cell therapy, which has potential in treating diseases such as multiple sclerosis. In this procedure, the patient’s immune system is suppressed and then re-established with the patient’s own stem cells, essentially ’resetting’ the immune system. Moreover, dietary intervention, such as a strict gluten-free diet, has been proven to manage celiac disease effectively. However, these treatments all have their own risks and side effects, and research is ongoing to refine these therapies and develop new ones.
Negative Passage Autoimmune disorders are a group of diseases where the body’s immune system attacks its own cells. There are many types of autoimmune diseases, including Rheumatoid Arthritis, Lupus, Celiac Disease, and Multiple Sclerosis. Each of these diseases has different symptoms, causes, and requires different treatments. Some common symptoms of autoimmune diseases are fatigue, joint pain, and swelling, skin problems, and abdominal pain. The causes of these diseases are not fully understood, but they are thought to be a combination of genetic and environmental factors. There is currently no cure for autoimmune diseases, but treatments can help manage the symptoms. Treatments include medication, physical therapy, and in some cases surgery. In the case of celiac disease, a strict gluten-free diet is necessary. It is important to work with a healthcare provider to develop a treatment plan that is tailored to the individual’s needs.
Table 7: Synthetic retrieval tasks and examples generated by GPT-4.
We present a list of generated retrieval scenarios as examples:
  • “Search for articles discussing the latest advancements in neurology.”
  • “Retrieval of articles discussing the symptoms and treatments of rare diseases given a query on rare diseases.”
  • “Find documents that discuss the impact of lifestyle changes on a specific medical condition.”
  • “Locate documents that provide information on the epidemiology of a certain disease in a specific region.”
[…]
Table 7 presents two illustrative examples where GPT-4 generates corresponding queries, positive passages, and negative passages for each synthetic retrieval task. The complete set of task names is provided in the supplementary materials.
Appendix B Task and Dataset Information
B.1 Pre-training Corpus
Dataset Size Line
PubMed Xiong et al. (2024) 8M https://huggingface.co/datasets/MedRAG/pubmed
arXiv, MedRxiv, BioRxiv Muennighoff et al. (2023) 577K https://huggingface.co/datasets/mteb/raw_arxiv
Meadow Wang et al. (2020) 460k https://huggingface.co/datasets/medalpaca/medical_meadow_cord19
Textbooks Jin et al. (2021) 50K https://huggingface.co/datasets/MedRAG/textbooks
StatPearls Xiong et al. (2024) 54K https://huggingface.co/datasets/MedRAG/statpearls
LitCovid Chen et al. (2021) 70K https://huggingface.co/datasets/KushT/LitCovid_BioCreative
S2ORC Lo et al. (2020) 600K https://github.com/allenai/s2orc
MS Marco Bajaj et al. (2016) 1.2M https://huggingface.co/datasets/Tevatron/msmarco-passage-corpus
Table 8: Biomedical corpora collection for unsupervised contrastive pre-training. : We randomly select 8M corpus from the full collections.
We publicly release the training recipe used in both the pre-training and fine-tuning stages to ensure transparency, reproducibility, and potential applicability to new domains. To equip BMRetriever with a strong foundation in biomedical contexts, we compile a diverse corpus of biomedical data sources. Table 8 summarizes the unlabeled corpora used for contrastive pre-training of our model, including their sizes and public availability. For pre-training on BMRetriever-7b, we only use 1M passages due to the efficiency issue.
For queries and passages, the instruction used in the contrastive pre-training stage is “Given a query, retrieve passages that are relevant to the query. Query: {}”, “Represent this passage. Passage: {}”.
B.2 Fine-tuning Task and Dataset
Dataset Size Task Link Instruction Format
BioMedical Domain
StackExchange Team (2021) 43K QA https://huggingface.co/datasets/flax-sentence-embeddings/stackexchange_titlebody_best_voted_answer_jsonl Given a biological query from the stackexchange, retrieve replies most relevant to the query
MedNLI Shivade (2017) 4.6K Sentence Similarity https://physionet.org/content/mednli/1.0.0/ Given a sentence, retrieve sentences with the same meaning
MQP McCreery et al. (2020) 3K Sentence Similarity https://huggingface.co/datasets/medical_questions_pairs Given a sentence, retrieve sentences with the same meaning
MedQuad Ben Abacha et al. (2019) 47K QA https://huggingface.co/datasets/lavita/MedQuAD Given a question, retrieve relevant documents that answer the question
HealthcareMagic Li et al. (2023b) 30K Dialogue https://huggingface.co/datasets/medical_dialog Given a question with context from online medical forums, retrieve responses that best answer the question
General Domain
ELI5 Fan et al. (2019) 18K Longform QA https://huggingface.co/datasets/eli5 Given a question, retrieve the highest voted answers on Reddit forum
GooAQ Khashabi et al. (2021) 100K QA https://huggingface.co/datasets/gooaq Given a question, retrieve relevant passages that answer the question
MS Marco Bajaj et al. (2016) 500K Web Search https://huggingface.co/datasets/ms_marco Given a web search query, retrieve relevant passages that answer the query
NQ Kwiatkowski et al. (2019) 58K QA https://github.com/facebookresearch/DPR/blob/main/dpr/data/download_data.py Given a question, retrieve Wikipedia passages that answer the question
FEVER Thorne et al. (2018) 10K Fact Verification https://huggingface.co/datasets/BeIR/fever Given a claim, retrieve documents that support or refute the claim
HotpotQA Yang et al. (2018) 8K Question Answering https://huggingface.co/datasets/BeIR/hotpotqa Given a complex question, retrieve documents that answer the question
NLI Bowman et al. (2015) 138K Natural Language Inference https://github.com/princeton-nlp/SimCSE/blob/main/data/download_nli.sh Given a premise, retrieve hypotheses that are entailed by the premise
Table 9: Labeled data collection for instruction fine-tuning with a diverse range of tasks, including both sentence-level NLI and passage-level QA. : Only a subset of the original dataset is sampled.
Real Datasets.
Table 9 displays the datasets used for instruction fine-tuning besides synthetic augmentation, which include a diverse range of tasks at both the sentence and passage levels across biomedical and general domains. Biomedical datasets cover biomedical QA (Team 2021, Ben Abacha et al. 2019), sentence similarity (Shivade 2017, McCreery et al. 2020), and dialogue (Li et al. 2023b). General domain datasets tackle long-form QA (Fan et al. 2019), web search (Bajaj et al., 2016), open-domain QA (Khashabi et al. 2021, Kwiatkowski et al. 2019), fact verification (Thorne et al. 2018), NLI (Bowman et al. 2015), and web search (Bajaj et al. 2016). For MS Marco [^5] https://msmarco.z22.web.core.windows.net/msmarcoranking/triples.train.small.tar.gz and NQ dataset [^6] https://dl.fbaipublicfiles.com/dpr/data/retriever/biencoder-nq-adv-hn-train.json.gz, we use the ground-truth annotations as well as the provided hard negative to form the fine-tuning data.
For non-retrieval tasks, we convert them into a retrieval format as follows:
  • For standard QA datasets, we directly use the question as the query and the gold evidence passages as the ground-truth passages.
  • For NLI and sentence similarity tasks, we treat sentence pairs identified as "entail" or "similar" as positive examples, while those labeled as "contradict" or "non-similar" serve as hard negatives.
  • For medical dialogue datasets, we consider the answer to the user query as the ground-truth passage for retrieval.
The query instructions are listed in the corresponding tables, while for passages, we use the same instruction format as the template used in the pre-training stage: “Represent this passage. Passage: {}”.
Synthetic Datasets.
We leverage LLM-generated synthetic data to augment the training set. For the query generation scenario, we generate 500K synthetic queries. After round-trip filtering, we retain approximately 420K (query, passage) pairs. The instructions used for generating synthetic queries are:
  • For the PubMed corpus: “Given a question, retrieve Pubmed passages that answer the question.
  • For the Meadow corpus on COVID-19: “Given a query on COVID-19, retrieve COVID-19 related articles that answer the query.
We generate 20,000 synthetic tasks and query-passage pairs using GPT-4. Table 7 presents some examples of synthetic retrieval tasks and query-passage pairs.
B.3 Evaluation Task and Dataset
Dataset Task # Queries # Documents Link Instruction Format
NFCorpus Boteva et al. (2016) Biomedical Search 323 3.6K https://huggingface.co/datasets/BeIR/nfcorpus Given a question, retrieve relevant documents that best answer the question
SciFact Wadden et al. (2020) Fact Verification 300 5K https://huggingface.co/datasets/BeIR/scifact Given a scientific claim, retrieve documents that support or refute the claim
SciDocs Cohan et al. (2020) Citation Prediction 1,000 25K https://huggingface.co/datasets/BeIR/scidocs Given a scientific paper title, retrieve paper abstracts that are cited by the given paper
Trec-COVID Voorhees et al. (2021) Biomedical Search 50 171K https://huggingface.co/datasets/BeIR/trec-covid Given a query on COVID-19, retrieve documents that answer the query
BIOSSES Soğancıoğlu et al. (2017) Biomedical Sentence Similarity 100 https://huggingface.co/datasets/biosses Given a sentence, retrieve sentences with the same meaning
BioASQ Tsatsaronis et al. (2015) Biomedical QA 500 500K http://participants-area.bioasq.org/datasets/ Given a question, retrieve Pubmed passages that answer the question
PubMedQA Jin et al. (2019) Biomedical QA 500 211K https://huggingface.co/datasets/qiaojin/PubMedQA Given a question, retrieve Pubmed passages that answer the question
iCliniq Chen et al. (2020) Biomedical CQA 7.3K 7.3K https://huggingface.co/datasets/medical_dialog Given a question with context from online medical forums, retrieve responses that best answer the question
DrugBank Wishart et al. (2018) Biomedical Entity Linking 4.1K 4.1K https://go.drugbank.com/ Given a drug, retrieve passages for its definition
MeSH Lipscomb (2000) Biomedical Entity Linking 29.6K 29.6K https://www.nlm.nih.gov/databases/download/mesh.html Given a concept, retrieve passages for its definition
RELISH Singh et al. (2023); Brown et al. (2019) Biomedical Paper Recommendation 3.2K 191.2K https://huggingface.co/datasets/allenai/scirepeval/viewer/relish Given an article, retrieve Pubmed articles that are relevant to this article
Table 10: Evaluation datasets for biomedical text representation tasks and retrieval-oriented downstream applications.
We conduct a comprehensive evaluation of BMRetriever on eleven datasets (Table 10) across five biomedical tasks, including:
Information Retrieval.
For passage retrieval tasks in biomedicine, we select four datasets from the BEIR benchmark (Thakur et al., 2021), each focusing on biomedical or scientific-related IR tasks involving complex, terminology-rich documents: (1) NFCorpus (Boteva et al., 2016) contains 323 queries related to nutrition facts for medical IR, sourced from 3.6K PubMed documents; (2) SciFact (Wadden et al., 2020) includes 300 queries, aiming to retrieve evidence-containing abstracts from 5K scientific papers for fact-checking; (3) SciDocs (Cohan et al., 2020) consists of 25K scientific papers for citation prediction with 1K queries containing article titles; (4) TREC-COVID (Voorhees et al., 2021) includes 50 queries, with an average of 493.5 relevant documents per query, specifically curated for biomedical IR related to COVID-19.
Sentence Similarity.
For sentence retrieval tasks, we evaluate retrieval models on (5) BIOSSES (Soğancıoğlu et al., 2017), which comprises 100 sentence pairs extracted from PubMed articles. The similarity of each sentence pair is annotated using a 5-point scale, ranging from 0 (no relation) to 4 (equivalent).
Question-and-Answering.
Besides passage and sentence retrieval tasks, we further evaluate the effectiveness of retrieval models on several retrieval-oriented downstream tasks, including biomedical QA. (6) BioASQ (Tsatsaronis et al., 2015) and (7) PubMedQA (Jin et al., 2019) are large-scale biomedical multi-choice QA datasets derived from PubMed articles. (8) iCliniq (Chen et al., 2020) contains medical QA pairs from the public health forum derived from conversations between clinicians and patients.
Entity Linking.
For additional retrieval-oriented downstream applications, we conduct two biomedical entity-linking experiments: (9) DrugBank (Wishart et al., 2018) for drug entity matching, and (10) MeSH (Lipscomb, 2000) for biomedical concept linking.
Paper Recommendation.
We evaluate the performance of retrieval models on a paper recommendation task using the (11) RELISH dataset (Singh et al., 2023; Brown et al., 2019). It assigns similarity scores ranging from 0 (not similar) to 2 (similar) for locating relevant literature from more than 180K PubMed abstracts.
Appendix C Baseline Information
We consider both sparse and dense retrieval models to provide a comprehensive evaluation of retrieval models in biomedical applications.
C.1 Baselines for Retrieval Tasks in Main Experiments
Sparse Retrieval Models.
Sparse retrieval models rely on lexical matching between query and document terms to calculate similarity scores.
  • BM25 (Robertson et al., 2009) is the most commonly used sparse retrieval model for lexical retrieval, employing a scoring function that calculates the similarity between two high-dimensional sparse vectors based on token matching and weighting.
Dense Retrieval Models.
Dense retrieval models utilize dense vector representations to capture semantic similarity between queries and documents. In our experiments, we consider dense retrieval models at various scales for a comprehensive evaluation: (1) Base Size (<1B parameters), (2) Large Size (1B-5B), and (3) XL Size (>5B).
  • Contriever (Izacard et al., 2022) is a dense retrieval model (110M) pre-trained via contrastive learning on documents sampled from Wikipedia and CC-Net (Wenzek et al., 2020) corpora.
  • Dragon (Lin et al., 2023) is a BERT-base-sized dense retrieval model (110M) that undergoes progressive training using a data augmentation approach, incorporating diverse queries and sources of supervision.
  • SPECTER 2.0 (Singh et al., 2023) is a scientific document representation model (110M) pretrained using multi-format representation learning, enabling tailored embeddings for a diverse range of task formats.
  • SciMult (Zhang et al., 2023) is a scientific dense retrieval model (110M) that employs a multi-task contrastive learning framework with task-aware specialization and instruction tuning to enhance performance on scientific literature retrieval tasks.
  • COCO-DR (Yu et al., 2022) is a lightweight dense retrieval model (110M) pre-trained using continuous contrastive learning and implicit distributionally robust optimization on domain-specific corpora, enabling adaptation to various downstream tasks, including those in the biomedical domain.
  • QExt (Meng et al., 2022) is a data augmentation method that trains dense retrieval models by selecting salient spans from the original document, and generating pseudo queries using transferred language models. We report the performance of QExt (110M) in the unsupervised dense retrieval setting only.
  • SGPT (Muennighoff, 2022) is a dense retrieval model that employs position-weighted mean pooling and fine-tunes only bias tensors to learn effective representations for semantic search. We conduct a comprehensive comparison with SGPT at different scales, including base size (125M), large size (1.3B and 2.7B), and XL size (5.8B).
  • MedCPT (Jin et al., 2023) is a biomedical embedding model (220M) specifically designed for biomedical literature retrieval, leveraging contrastive pre-training on medical corpora consisting of 255M user clicks from PubMed search logs (Fiorini et al., 2018).
  • GTR (Ni et al., 2022) is a generalizable dense retriever that initializes its dual encoders from T5 (Raffel et al., 2020). GTR is pre-trained on a private Community QA dataset and fine-tuned using the NQ (Kwiatkowski et al., 2019) and MS MARCO (Bajaj et al., 2016) datasets. We conduct a comprehensive comparison with GTR at varying scales, including GTR-Large (335M), GTR-XL (1.2B), and GTR-XXL (4.8B).
  • InstructOR (Su et al., 2023) is a multitask embedder that generates task- and domain-aware embeddings for a given text input and its corresponding task instructions, without requiring any additional training. We evaluate InstructOR at both base (335M) and large (1.5B) scales.
  • E5-Large-v2 (Wang et al., 2022b) adopts a complex multi-stage training paradigm that first pre-trains on large-scale weakly-supervised text pairs and then fine-tunes on several labeled datasets. Due to the knowledge distillation from a teacher model, which involves additional supervised training signals, we consider E5-Large-v2 (335M) as an unfair comparison.
  • BGE-Large (Chen et al., 2024) is a dense retrieval model (335M) that uses graph-based embedding techniques and a multi-stage training paradigm similar to E5 (Wang et al., 2022b). However, since BGE employs a hybrid retrieval strategy incorporating additional lexical and multi-vector retrieval, we consider it an unfair comparison for our focus on dense retrieval only. We acknowledge BGE as a concurrent work and provide its performance for reference purposes.
  • LLaRA (Li et al., 2023a) uses LLM-generated text embeddings to reconstruct input sentence tokens and predict next sentence tokens.
  • RepLLaMA (Ma et al., 2023) is a dense retriever (7B) that fine-tunes the LLaMA model for effective representation learning in passage and document retrieval using the MS MARCO datasets (Bajaj et al., 2016).
  • LLM2Vec (BehnamGhader et al., 2024) is an unsupervised approach that transforms LLMs into text encoders by enabling bidirectional attention via masked next token prediction and adopts unsupervised contrastive learning for sequence representation learning.
  • Grit-LM (Muennighoff et al., 2024) is a generative dense retrieval model (7B) that incorporates generative representational instruction tuning, enabling the model to handle both generative and embedding tasks by distinguishing between them through instructions.
  • E5-Mistral (Wang et al., 2024b) is an enhanced version of the E5 (Wang et al., 2022b) that incorporates synthetic data generated by LLMs for a diverse range of text embedding tasks.
  • CPT-text (Neelakantan et al., 2022) is a dense retrieval model pre-trained on web-scale data using a contrastive objective with neighboring text passages as positive pairs. We only consider its performance as a reference rather than a fair comparison due to its large size (175B parameters).
C.2 Baselines for Retrieval-Oriented Downstream Applications
In experiments for retrieval-oriented downstream applications, we only compare BMRetriever to the strongest, most relevant, and fair baselines, including: (1) Base Size (<1B): Dragon (Lin et al., 2023), MedCPT (Jin et al., 2023), and E5-Large-v2 (Wang et al., 2022b); (2) Large Size (1B-5B): InstructOR (Su et al., 2023) and SGPT-2.7B (Muennighoff, 2022); and (3) XL Size (>5B): E5-Mistral (Wang et al., 2024b).
Appendix D Implementation Details
The backbones used for BMRetriever are available in Table 1. The learning rates are set to $55e-$ 5 for the 410M and 1B variants, $54e-$ 5 for the 2B variant, and $52e-$ 5 for the 7B variant during pre-training; $55e-$ 5 for the 410M and 1B variants, $52e-$ 5 for the 2B variant, and $51e-$ 5 for the 7B variant during fine-tuning. The global batch size is set to $256$ for the 410M and 1B variants, $128$ for the 2B variant, and $64$ for 7B variants. To optimize GPU memory consumption, we train our models with LoRA ($r=16$, $\alpha=32$(Hu et al., 2022), brain floating point (bfloat16) quantization, and DeepSpeed gradient checkpointing (Rasley et al., 2020). The training is performed on 4 NVIDIA H100 GPUs for 2 epochs during pre-training and 1 epoch during fine-tuning, using a maximum sequence length of 512 tokens. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with a linear learning rate warm-up for the first 100 steps. For contrastive learning, we set $\tau=1$ without any further tuning.
Appendix E Cosine Similarity v.s. Dot Product
We explore different objectives for embedding similarity, namely dot product and cosine similarity. From the experimental results in Figure 5, we empirically observe that the dot product could achieve a better empirical performance. Thus, we choose to use dot product by default as our similarity metrics.
Figure 5: Comparison of performance using dot product and cosine similarity.
Appendix F Similarity Score
Figure 6 depicts the distributions of cosine similarity scores for positive and negative embedding pairs across two datasets. The left side displays the similarity distributions for negative examples, while the right side shows the distributions for positive examples. These figures illustrate that BMRetriever exhibits a larger separation between positive and negative examples, showing its enhanced ability to effectively retrieve relevant passages.
(a) SciFact
(b) iCliniq
Figure 6: The cosine similarity on positive pair embeddings and negative pair embeddings.
Appendix G Efficiency
Table 11 exhibits the document encoding speed and retrieval latency of BMRetriever and baseline dense retrieval models. While BMRetriever introduces additional encoding latency compared to BERT-based retrievers, we do not incorporate significant overhead when compared to baselines of similar model size.
Models Size Document Encoding Speed (# docs / s / GPU) Retrieval Latency (ms)
MedCPT Jin et al. (2023) 220M 1390.1 11.6
InstructOR Su et al. (2023) 1.5B 181.2 14.6
SGPT Muennighoff (2022) 2.7B 98.5 35.5
E5-Mistral Wang et al. (2024b) 7B 51.8 58.6
BMRetriever 410M 471.2 14.6
BMRetriever 1B 194.0 28.6
BMRetriever 2B 166.2 28.6
BMRetriever 7B 51.8 58.6
Table 11: Time complexity of BMRetriever.